OpenStudy (anonymous):

two cars leave town traveling in opposite directions. one travels at 80km/hr and the other at 96 km/h. in how many hours will they be 528 km aprt? )TWO VARIABLES ???? O_________o)

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

OpenStudy (perl):

Use parametric equations to model their motion x = 80*t x = -96 * t

OpenStudy (perl):

now the distance between x1 and x2 is 80*t - (-96t)

OpenStudy (perl):

so solve 176t = 528

OpenStudy (perl):

@kropot72 what were you going to type? sorry

OpenStudy (kropot72):

@perl Basically what you said. np.

OpenStudy (perl):

thanks, you probably can draw it better (i dont use latex :/ )

OpenStudy (anonymous):

woah, hopefully i dont need to deal with this on my quiz...